OpenStudy (anonymous):

If M is the product of all integers from 1 to 50 which a factor 2^n...then the maximum value of n is ?????

OpenStudy (anonymous):

is it 47

OpenStudy (anonymous):

use greatest integer function for all and then add [50/2] +[50/4]+[50/8]+[50/16]+[50/32]=25+12+6+3+1

OpenStudy (anonymous):

@shubhamsrg going thorugh NEST2010

OpenStudy (anonymous):

yeah xactly! u toooo givin nest! ?

OpenStudy (anonymous):

no dear...

OpenStudy (anonymous):

so u do u knw?

OpenStudy (anonymous):

xams are near thats why asked

OpenStudy (anonymous):

good lust/love for maths is mportant for these kind of xams best of luck....

OpenStudy (shubhamsrg):

well,, 1*2*3*4...50 = 1*2*3*(2*2)*....25*(13*2)*27*(14*2)*29*(15*2)....49*(25*2) =(1*2*3*4...25)^2* (2^25 )..solving the inner bracket,,you get 2^12 outside..doing again,,you get 2^6 outside..you finally get what @matricked said!!

OpenStudy (anonymous):

no not like that u get the abv rules in most of the text books

OpenStudy (shubhamsrg):

this was just another method,,am sure there must be many such,..